OpenStudy (anonymous):

Jon and Becky measured the circle-shaped part of a sun they drew on the sidewalk. The diameter of the circle is 59 centimeters. What is the approximate area of the circle? (Use 3.14 for .) 10,930.34 cm2 2,732.585 cm2 683.14625 cm2 185.26 cm2

OpenStudy (aaronandyson):

Area = pi*r^2 Diameter/2 = r 59/2 = r

OpenStudy (anonymous):

can u explain a lil more please...? im sorry im terrible at math

OpenStudy (aaronandyson):

LOL I was just like you but take it easy can you'll be a master at it !:P

OpenStudy (aaronandyson):

Moving back to the question Let's write down all the information we have.

OpenStudy (anonymous):

ok

OpenStudy (aaronandyson):

We have been given that the diameter of this circle is 59 cms

OpenStudy (anonymous):

whats diameter i forgot LoL

OpenStudy (aaronandyson):

When we have the diameter divide it by 2

OpenStudy (aaronandyson):



OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

and divide it by 2?

OpenStudy (aaronandyson):

Yes.

OpenStudy (anonymous):

29.5 ???

OpenStudy (aaronandyson):

Yes!

OpenStudy (aaronandyson):

So,29.5cms is the radius of the circle.

OpenStudy (aaronandyson):

Agreed?

OpenStudy (anonymous):

is that it???

OpenStudy (aaronandyson):

Nope Area of circle = pi*radius Area of circle = 3.14*29.5

OpenStudy (anonymous):

ok whats radius...? im sorry i didnt really do classes this year and last year i didnt listen cause my math teacher was being mean

OpenStudy (aaronandyson):